# Dynamic Population Games: A Tractable Intersection of Mean-Field Games and Population Games

### v2

@@ Abstract
In many real-world large-scale decision problems, self-interested agents have individual dynamics and optimize their own long-term payoffs. Important examples include the competitive access to shared resources (e.g., roads, energy, or bandwidth) but also non-engineering domains like epidemic propagation and control. These problems are natural to model as mean-field games. However, existing mathematical formulations of mean field games have had limited applicability in practice, since they require solving non-standard initial-terminal-value problems that are tractable only in limited special cases. In this letter, we propose a novel formulation, along with computational tools, for a practically relevant class of Dynamic Population Games (DPGs), which correspond to discrete-time, finite-state-and-action, stationary mean-field games. Our main contribution is a mathematical reduction of Stationary Nash Equilibria (SNE) in DPG s to standard Nash Equilibria (NE) in static population games. This reduction is leveraged to guarantee the existence of a SNE, develop an evolutionary dynamics-based SNE computation algorithm, and derive simple conditions that guarantee stability and uniqueness of the SNE. Additionally, DPG s enable us to tractably incorporate multiple agent types, which is of particular importance to assess fairness concerns in resource allocation problems. We demonstrate our results by computing the SNE in two complex application examples: fair resource allocation with heterogeneous agents and control of epidemic propagation.
@@
I Introduction
@@
In all of these applications, the availability of practical computation tools to predict the Mean Field Equilibrium (MFE) is instrumental to study the effect of complex control or policy design choices. However, existing formulations of mean field games face a computational bottleneck as they require solving two coupled partial differential equations (PDEs): a Hamilton-Jacobi-Bellman (HJB) equation characterizing the optimality of the agents’ policy; and a Fokker-Planck-Kolmogorov (FPK) equation ensuring that the mean field is consistent with the state dynamics under the optimal policy. The resulting initial-terminal-value problem is highly non-standard, and computation tools are not yet fully mature, since the existence (and uniqueness) of solutions is guaranteed under abstract assumptions that are difficult to verify in practice [1, 2, 14, 15, 16, 17]. In practice, these assumptions typically manifest in linear dynamics with quadratic costs [8, 9, 10, 18], and/or a weak dependency on the mean field (e.g., through the average state or action only [8, 10, 19]).
@@
On the other hand, the well-understood class of static population games provides a rich set of Nash Equilibrium (NE) analysis and computation tools [20]. In population games, a large population of agents optimize immediate payoffs over a finite set of actions, and the agents do not have individual state dynamics. The macroscopic evolution of actions is modelled by evolutionary dynamics, which include the well-known replicator, best response, Smith, and projection dynamics [20]. Evolutionary dynamics are useful for NE computation as their stationary points coincide with NE, and they are guaranteed to converge in the well-studied class of stable population games [20]. Moreover, recent works have enriched the classical evolutionary dynamics with passivity-based and higher-order learning techniques that extend the classes of games in which convergence is guaranteed [21, 22, 23, 24, 25].
@@
Population games are trivially mean field games, in which agents have only one state (with trivial dynamics) and finite actions. The main contribution of this letter is to show that there is a larger and meaningful intersection: many practically relevant mean field games are population games. Those are the class of discrete-time, finite-state-and-action, stationary mean-field games, which we call Dynamic Population Games (DPGs) for short. A natural solution concept in this class of games is the Stationary Nash Equilibrium (SNE) [4, 26], also referred to as the stationary mean field equilibrium [5, 27, 28], in which microscopically, each agent follows an optimal policy, and macroscopically, the state distribution is stationary. We show a mathematical reduction of SNE to standard NE in population games: For each DPG, there exists a suitably defined static population game whose NE coincide with SNE of the DPG.
@@
The reduction of SNE in DPG s to NE in population games unlocks a rich set of tools for SNE analysis and computation. We leverage it to guarantee the existence of a SNE, develop an evolutionary dynamics-based SNE computation algorithm, and provide simple conditions that guarantee stability and uniqueness of the SNE. Finally, the versatility of our approach is demonstrated in two complex application examples involving fair resource allocation to heterogeneous agents, as well as epidemic propagation and control. Open source software for SNE computation is provided in: https://gitlab.ethz.ch/elokdae/dynamic-population-games.
@@
The remainder of the letter is organized as follows: Section II introduces notation and definitions for DPG s; Section III presents our main result regarding the reduction of SNE to NE in population games; Section IV discusses important consequences of the reduction; and Section V presents numerical results from two application examples. The letter is concluded in Section VI.
I-A Notation
@@
Let $a,d\in D\subseteq{\mathbb{N}}$ and let $c\in C\subseteq{\mathbb{R}}^{n}$, then for a function $f:D\times C\rightarrow{\mathbb{R}}$, we distinguish discrete and continuous arguments through the notation $f[d](c)$. Alternatively, we write $f:C\rightarrow{\mathbb{R}}^{|D|}$ as the vector-valued function $f(c)$, with $f[d](c)$ denoting its $d$-th element. Similarly, $p[a\mid d](c)$ denotes the conditional probability of $a$ given $d$ and $c$. Specifically, $p[d^{+}\mid d](c)$ denotes one-step transition probabilities. For $m\in{\mathbb{R}}_{+}$ denoting a measure of total mass, we denote by $\delta\in m\>\Delta(D):=\left\{\left.\sigma\in{\mathbb{R}}_{+}^{|D|}\right\rvert\sum_{d\in D}\sigma[d]=m\right\}$ a distribution over the elements of $D$, with $\delta[d]\in[0,m]$ denoting the mass of element $d$ (if $m=1$ then $\delta$ is a probability distribution). Finally, when considering heterogeneous agent types, we denote by $x_{\tau}$ a quantity associated to type $\tau$.
II Dynamic Population Games (DPGs)
@@
II-A Dynamic Population Game (DPG)Setting
@@
At discrete time steps, agents participate in a game in which each agent of type-state $[\tau,x]$ plays one of a finite number of actions $a\in{\mathcal{A}}_{\tau}[x]$. Type $\tau$ agents pick their action according to the stationary homogeneous policy $\pi_{\tau}:{\mathcal{X}}\rightarrow\Delta({\mathcal{A}}_{\tau}[x])$, which maps the state to a probability distribution over the actions. The probability to play action $a$ when in state $x$ is denoted by $\pi_{\tau}[a\mid x]\in[0,1]$. The set of policies of type $\tau$ is denoted by $\Pi_{\tau}$, and the concatenation of the policies is denoted by $\pi=(\pi_{1},\dots,\pi_{N_{\tau}})\in\Pi=\prod_{\tau\in{\mathcal{T}}}\Pi_{\tau}$.
The pair $(d,\pi)\in{\mathcal{D}}\times\Pi$ is referred to as either the mean field (in the tradition of mean field games), or the social state (in the tradition of population games; a formal equivalency between mean fields and social states is shown in Section III), since $(d,\pi)$ gives a macroscopic description of the distribution of the agents’ types and states, as well as how they behave.
@@
II-B Solution Concept: Stationary Nash Equilibrium (SNE)
Before introducing the solution concept in DPG s, we must introduce definitions pertaining to the strategic problem faced by the population agents. Each agent faces a discounted Markov Decision Process (MDP), i.e., it aims to maximize its discounted infinite horizon payoff, with the type-dependant discount factor $\alpha_{\tau}\in[0,1)$. The individual MDP of each agent is coupled to other agents through the mean field $(d,\pi)$. Namely, when a type $\tau$ agent follows policy $\pi_{\tau}$, its expected immediate payoff and state transition matrix are respectively given by,
@@ II-B Solution Concept: Stationary Nash Equilibrium (SNE)
$Q_{\tau}[x,a](d,\pi)\\ =r_{\tau}[x,a](d,\pi)+\alpha_{\tau}\sum_{x^{+}\in{\mathcal{X}}}p_{\tau}[x^{+}\mid x,a](d,\pi)\>V_{\tau}[x^{+}](d,\pi).$ (6)
Equation (6) is the expected infinite horizon payoff if the type $\tau$ agent in state $x$ deviates from $\pi_{\tau}$ for a single stage by playing $a$, then follows $\pi_{\tau}$ in the future.
It is well-known that a policy attains the maximum value function (5) if and only if it maximizes the state-action value function (6) for all states $x\in{\mathcal{X}}$ [29, Proposition 1.3.4]. This motivates the definition of the state-dependent best response correspondence
@@
III Main Result: Reduction to Population Game
@@
Proof:
The proof proceeds by construction. We construct a population game $F^{\mathcal{G}}$ with $N_{\rho}=N_{\tau}\>(1+N_{x})$ populations. The the first $N_{\tau}$ populations correspond to the different types, and we denote these populations as $\rho=\tau\in{\mathcal{T}}$. The remaining $N_{\tau}N_{x}$ populations correspond to the type-state pairs, and we denote these populations as $\rho=[\tau,x]\in{\mathcal{T}}\times{\mathcal{X}}$. The action space of populations $\rho=\tau$ is the state space ${\mathcal{X}}$, i.e., ${\mathcal{A}}_{\tau}={\mathcal{X}}$, whereas the action space of populations $\rho=[\tau,x]$ is respectively ${\mathcal{A}}_{[\tau,x]}={\mathcal{A}}_{\tau}[x]$. The mass of populations $\rho=\tau$ is respectively $g_{\tau}$, whereas the mass of populations $\rho=[\tau,x]$ is uniformly $g_{[\tau,x]}=1$. With this construction, it holds that ${\mathcal{S}}={\mathcal{D}}\times\Pi$, i.e., each social state $s\in{\mathcal{S}}$ under $F^{\mathcal{G}}$ is formally equivalent to a mean field $(d,\pi)\in{\mathcal{D}}\times\Pi$ under ${\mathcal{G}}$. We thus write $s=(d,\pi)$ and define the payoffs $F^{\mathcal{G}}(d,\pi)$ for both groups of populations as
@@
To prove the theorem, we now show that the NE condition of $F^{\mathcal{G}}$, i.e., $s^{*}_{\rho}\in B_{\rho}(s^{*})$ for all $\rho\in{\mathcal{P}}$, is formally equivalent to conditions (SNE.1)–(SNE.2), with $s^{*}=(d^{*},\pi^{*})$. By construction, for all populations $\rho=[\tau,x]\in{\mathcal{T}}\times{\mathcal{X}}$, with $s^{*}_{[\tau,x]}=\pi^{*}_{\tau}[\cdot\mid x]$, condition $\pi^{*}_{\tau}[\cdot\mid x]\in B_{[\tau,x]}(d^{*},\pi^{*})$ is formally equivalent to condition (SNE.2). Therefore, we must show that for the remaining populations $\rho=\tau\in{\mathcal{T}}$, with $s^{*}_{\tau}=d^{*}_{\tau}$, $d^{*}_{\tau}\in B_{\tau}(d^{*},\pi^{*})$ if and only if $d^{*}_{\tau}[x]$ satisfies condition (SNE.1) for all $x\in{\mathcal{X}}$. One direction of the implication is trivial: if $d^{*}_{\tau}[x]$ satisfies condition (SNE.1) for all $[\tau,x]\in{\mathcal{T}}\times{\mathcal{X}}$, then $F^{\mathcal{G}}_{\tau}[x](d^{*},\pi^{*})=0$, and all state distributions satisfy $d_{\tau}\in B_{\tau}(d^{*},\pi^{*})$, including $d^{*}_{\tau}$. For the other direction of the implication, suppose, for the sake of contradiction, that $d^{*}_{\tau}\in B_{\tau}(d^{*},\pi^{*})$ for all $\tau\in{\mathcal{T}}$, but there is some type-state $[\tilde{\tau},\tilde{x}]\in{\mathcal{T}}\times{\mathcal{X}}$ for which $F^{\mathcal{G}}_{\tilde{\tau}}[\tilde{x}](d^{*},\pi^{*})\neq 0$. Notice that $\sum_{x\in{\mathcal{X}}}F^{\mathcal{G}}_{\tau}[x](d,\pi)=0$ for all $\tau\in{\mathcal{T}}$ and $(d,\pi)\in{\mathcal{D}}\times{\mathcal{X}}$; therefore, if $F^{\mathcal{G}}_{\tilde{\tau}}[\tilde{x}](d^{*},\pi^{*})>0$ (respectively, $<0$), there must be another state $\hat{x}\in{\mathcal{X}}$ for which $F^{\mathcal{G}}_{\tilde{\tau}}[\hat{x}](d^{*},\pi^{*})<0$ (respectively, $>0$). Suppose without loss of generality that $F^{\mathcal{G}}_{\tilde{\tau}}[\tilde{x}](d^{*},\pi^{*})>0$ and $F^{\mathcal{G}}_{\tilde{\tau}}[\hat{x}](d^{*},\pi^{*})<0$. Since $d^{*}_{\tilde{\tau}}\in B_{\tilde{\tau}}(d^{*},\pi^{*})$, it must be that $d^{*}_{\tilde{\tau}}[\hat{x}]=0$. But this contradicts $F^{\mathcal{G}}_{\tilde{\tau}}[\hat{x}](d^{*},\pi^{*})<0$, in which $-d^{*}_{\tilde{\tau}}[\hat{x}]$ is the only negative term. We thus showed that if $d^{*}_{\tau}\in B_{\tau}(d^{*},\pi^{*})$ for all $\tau\in{\mathcal{T}}$, it must be that $F^{\mathcal{G}}_{\tau}[x](d^{*},\pi^{*})=0$ for all $[\tau,x]\in{\mathcal{T}}\times{\mathcal{X}}$, i.e., $d^{*}_{\tau}$ satisfies condition (SNE.1) for all $[\tau,x]\in{\mathcal{T}}\times{\mathcal{X}}$. ∎
IV Consequences of the Reduction

### v3

# {textblock*}

@@ Abstract
In many real-world large-scale decision problems, self-interested agents have individual dynamics and optimize their own long-term payoffs. Important examples include the competitive access to shared resources (e.g., roads, energy, or bandwidth) but also non-engineering domains like epidemic propagation and control. These problems are natural to model as mean-field games. Existing mathematical formulations of mean field games have had limited applicability in practice, since they require solving non-standard initial-terminal-value problems that are tractable only in limited special cases. In this letter, we propose a novel formulation, along with computational tools, for a practically relevant class of Dynamic Population Games (DPGs), which correspond to discrete-time, finite-state-and-action, stationary mean-field games. Our main contribution is a mathematical reduction of Stationary Nash Equilibria (SNE) in DPGs to standard Nash Equilibria (NE) in static population games. This reduction is leveraged to guarantee the existence of a SNE, develop an evolutionary dynamics-based SNE computation algorithm, and derive simple conditions that guarantee stability and uniqueness of the SNE. We provide two examples of applications: fair resource allocation with heterogeneous agents and control of epidemic propagation.
@@
{IEEEkeywords}
Game theory, Markov processes, Mean field games
1 Introduction
[…]
@@ 1 Introduction
In all of these applications, the availability of practical computation tools to predict the Mean Field Equilibrium (MFE) is instrumental to study the effect of complex control or policy design choices. However, general formulations of mean field games face a computational bottleneck as they require solving an initial-terminal-value problem consisting of a backward Hamilton-Jacobi-Bellman (HJB) equation characterizing the optimality of the agents’ policy trajectory; and a forward Fokker-Planck-Kolmogorov (FPK) equation ensuring that the mean field trajectory is consistent with the state dynamics under the optimal policy trajectory. Existence (and uniqueness) of solutions to the initial-terminal-value problem is guaranteed under abstract assumptions that are difficult to verify in practice [1, 2, 14, 15, 16, 17]. These assumptions typically manifest in linear dynamics with quadratic costs [8, 9, 10, 18], and/or a weak dependency on the mean field (e.g., through the average state or action only [8, 10, 19, 20]).
Due to these challenges, previous works have considered the less general solution concept of the stationary mean field equilibrium [5, 21, 22], also referred to as the Stationary Nash Equilibrium (SNE) [4, 23], which is a single pair of optimal policy and the associated stationary mean field. For the practically relevant class of discrete-time, finite-state-and-action, stationary mean-field games, which we call Dynamic Population Games (DPGs) for short, SNE are computationally attractive as they are characterized as finite-dimensional fixed points rather than initial-terminal-value problems. SNE computation has been the subject of recent works, and thus far computational tools are based on standard Picard-type fixed point methods [5, 22, 24].
@@
A Nash Equilibrium (NE) of a static population game is also characterized by a fixed point; however, there exists a rich set of NE analysis and computation tools beyond conventional fixed point methods [25]. In particular, evolutionary dynamics, which include the well-known replicator, best response, Smith, and projection dynamics [25], are useful for NE computation, since they leverage the game structure to ensure that stationary points coincide with NE, and the stability of these dynamics is well-understood [25]. Moreover, recent works have enriched the classical evolutionary dynamics with passivity-based and higher-order learning techniques that extend the classes of games in which convergence to a NE is guaranteed [26, 27, 28, 29, 30].
@@
Population games are trivially mean field games, in which agents have only one state (with trivial dynamics) and finite actions. The main contribution of this letter is to show that there is a larger and meaningful intersection: Dynamic Population Games (DPGs)are population games. We show a mathematical reduction of SNE to standard NE in population games: For each DPG, there exists a suitably defined static population game whose NE coincide with SNE of the DPG.
@@
The reduction of SNE in DPGs to NE in population games unlocks a rich set of tools for SNE analysis and computation. We leverage it to guarantee the existence of a SNE, develop an evolutionary dynamics-based SNE computation algorithm, and provide simple conditions that guarantee stability and uniqueness of the SNE. Finally, the versatility of our approach is demonstrated in two complex application examples involving fair resource allocation to heterogeneous agents, as well as epidemic propagation and control. Open source software for SNE computation is provided at:
https://gitlab.ethz.ch/elokdae/dynamic-population-games.
@@
The letter is organized as follows: Section 2 defines DPGs; Section 3 presents our main result: the reduction of SNE to NE in population games; Section 4 discusses important consequences of the reduction; and Section 5 presents two application examples. The letter is concluded in Section 6.
1.1 Notation
@@
Let $a,d\in D\subseteq{\mathbb{N}}$ and let $c\in C\subseteq{\mathbb{R}}^{n}$, then for a function $f:D\times C\rightarrow{\mathbb{R}}$, we distinguish discrete and continuous arguments through the notation $f[d](c)$. Alternatively, we write $f:C\rightarrow{\mathbb{R}}^{|D|}$ as the vector-valued function $f(c)$, with $f[d](c)$ denoting its $d$-th element. Similarly, $p[a\mid d](c)$ denotes the conditional probability of $a$ given $d$ and $c$. For $m\in{\mathbb{R}}_{+}$ denoting a measure of total mass, we denote by $\delta\in m\>\Delta(D):=\left\{\left.\sigma\in{\mathbb{R}}_{+}^{|D|}\right\rvert\sum_{d\in D}\sigma[d]=m\right\}$ a distribution over the elements of $D$, with $\delta[d]\in[0,m]$ denoting the mass of element $d$ (if $m=1$ then $\delta$ is a probability distribution). Finally, when considering heterogeneous agent types, we denote by $x_{\tau}$ a quantity associated to type $\tau$.
2 Dynamic Population Games (DPGs)
@@
2.1 Dynamic Population Game (DPG)Setting
@@
At discrete time steps, agents participate in a game in which each agent of type-state $[\tau,x]$ plays one of a finite number of actions $a\in{\mathcal{A}}_{\tau}[x]$. Type $\tau$ agents pick their action according to the stationary homogeneous policy $\pi_{\tau}:{\mathcal{X}}\rightarrow\Delta({\mathcal{A}}_{\tau}[x])$, which maps the state to a probability distribution over the actions [^1] This class of policies has been previously referred to as oblivious policies [4], since the agents choose their actions based on their individual state only and not the state distribution of others.. The probability to play action $a$ when in state $x$ is denoted by $\pi_{\tau}[a\mid x]\in[0,1]$. The set of policies of type $\tau$ is denoted by $\Pi_{\tau}$, and the concatenation of the policies is denoted by $\pi=(\pi_{1},\dots,\pi_{N_{\tau}})\in\Pi=\prod_{\tau\in{\mathcal{T}}}\Pi_{\tau}$.
The pair $(d,\pi)\in{\mathcal{D}}\times\Pi$ is referred to as the mean field (in the tradition of mean field games) or the social state (in the tradition of population games; a formal equivalency is shown in Section 3). It gives a macroscopic description of the distribution of the agents’ types and states, as well as how they behave [^2] Our definition of mean field $(d,\pi)$ differs slightly from the definition as the joint state-action distribution [3, 4, 5]; however, the latter can be readily derived from $(d,\pi)$..
@@
2.2 Solution Concept: Stationary Nash Equilibrium (SNE)
Before introducing the solution concept in DPG s, we must introduce definitions pertaining to the strategic problem faced by the population agents. Each agent faces a discounted Markov Decision Process (MDP), i.e., aims to maximize its discounted infinite horizon payoff, with the type-dependant discount factor $\alpha_{\tau}\in[0,1)$. Each agent’s MDP is coupled to other agents through the mean field $(d,\pi)$. Namely, when a type $\tau$ agent follows policy $\pi_{\tau}$, its expected immediate payoff and state transition matrix are respectively given by,
@@ 2.2 Solution Concept: Stationary Nash Equilibrium (SNE)
$Q_{\tau}[x,a](d,\pi)=r_{\tau}[x,a]+\alpha_{\tau}\sum_{x^{+}\in{\mathcal{X}}}p_{\tau}[x^{+}\mid x,a]\>V_{\tau}[x^{+}].$ (6)
Equation (6) is the expected infinite horizon payoff if the type $\tau$ agent in state $x$ deviates from $\pi_{\tau}$ for a single stage by playing $a$, then follows $\pi_{\tau}$ in the future. Notice that we omitted the dependency on $(d,\pi)$ on the right-hand side of Equations (5)–(6) for notational convenience.
A policy attains the maximum value function (5) if and only if it maximizes the state-action value function (6) for all states $x\in{\mathcal{X}}$ [31, Proposition 1.3.4]. This motivates the definition of the state-dependent best response correspondence
@@
3 Main Result: Reduction to Population Game
@@
Proof 3.2.
The proof proceeds by construction. We construct a population game $F^{\mathcal{G}}$ with $N_{\rho}=N_{\tau}\>(1+N_{x})$ populations. The first $N_{\tau}$ populations correspond to the different types, and we denote these populations as $\rho=\tau\in{\mathcal{T}}$. The remaining $N_{\tau}N_{x}$ populations correspond to the type-state pairs, and we denote these populations as $\rho=[\tau,x]\in{\mathcal{T}}\times{\mathcal{X}}$. The action space of populations $\rho=\tau$ is the state space ${\mathcal{X}}$, i.e., ${\mathcal{A}}_{\tau}={\mathcal{X}}$, whereas the action space of populations $\rho=[\tau,x]$ is respectively ${\mathcal{A}}_{[\tau,x]}={\mathcal{A}}_{\tau}[x]$. The mass of populations $\rho=\tau$ is respectively $g_{\tau}$, whereas the mass of populations $\rho=[\tau,x]$ is uniformly $g_{[\tau,x]}=1$. With this construction, it holds that ${\mathcal{S}}={\mathcal{D}}\times\Pi$, i.e., each social state $s\in{\mathcal{S}}$ under $F^{\mathcal{G}}$ is formally equivalent to a mean field $(d,\pi)\in{\mathcal{D}}\times\Pi$ under ${\mathcal{G}}$. We thus write $s=(d,\pi)$ and define the payoffs $F^{\mathcal{G}}(d,\pi)$ for both groups of populations as
@@
To prove the theorem, we now show that the NE condition of $F^{\mathcal{G}}$, i.e., $s^{*}_{\rho}\in B_{\rho}(s^{*})$ for all $\rho\in{\mathcal{P}}$, is formally equivalent to conditions (SNE.1)–(SNE.2), with $s^{*}=(d^{*},\pi^{*})$. By construction, for all populations $\rho=[\tau,x]\in{\mathcal{T}}\times{\mathcal{X}}$, with $s^{*}_{[\tau,x]}=\pi^{*}_{\tau}[\cdot\mid x]$, condition $\pi^{*}_{\tau}[\cdot\mid x]\in B_{[\tau,x]}(d^{*},\pi^{*})$ is formally equivalent to condition (SNE.2). Therefore, we must show that for the remaining populations $\rho=\tau\in{\mathcal{T}}$, with $s^{*}_{\tau}=d^{*}_{\tau}$, $d^{*}_{\tau}\in B_{\tau}(d^{*},\pi^{*})$ if and only if $d^{*}_{\tau}[x]$ satisfies condition (SNE.1) for all $x\in{\mathcal{X}}$. One direction of the implication is trivial: if $d^{*}_{\tau}[x]$ satisfies condition (SNE.1) for all $[\tau,x]\in{\mathcal{T}}\times{\mathcal{X}}$, then $F^{\mathcal{G}}_{\tau}[x](d^{*},\pi^{*})=0$, and all state distributions satisfy $d_{\tau}\in B_{\tau}(d^{*},\pi^{*})$, including $d^{*}_{\tau}$. For the other direction of the implication, suppose, for the sake of contradiction, that $d^{*}_{\tau}\in B_{\tau}(d^{*},\pi^{*})$ for all $\tau\in{\mathcal{T}}$, but there is some type-state $[\tilde{\tau},\tilde{x}]\in{\mathcal{T}}\times{\mathcal{X}}$ for which $F^{\mathcal{G}}_{\tilde{\tau}}[\tilde{x}](d^{*},\pi^{*})\neq 0$. Notice that $\sum_{x\in{\mathcal{X}}}F^{\mathcal{G}}_{\tau}[x](d,\pi)=0$ for all $\tau\in{\mathcal{T}}$ and $(d,\pi)\in{\mathcal{D}}\times{\mathcal{X}}$; therefore, if $F^{\mathcal{G}}_{\tilde{\tau}}[\tilde{x}](d^{*},\pi^{*})>0$ (respectively, $<0$), there must be another state $\hat{x}\in{\mathcal{X}}$ for which $F^{\mathcal{G}}_{\tilde{\tau}}[\hat{x}](d^{*},\pi^{*})<0$ (respectively, $>0$). Suppose without loss of generality that $F^{\mathcal{G}}_{\tilde{\tau}}[\tilde{x}](d^{*},\pi^{*})>0$ and $F^{\mathcal{G}}_{\tilde{\tau}}[\hat{x}](d^{*},\pi^{*})<0$. Since $d^{*}_{\tilde{\tau}}\in B_{\tilde{\tau}}(d^{*},\pi^{*})$, it must be that $d^{*}_{\tilde{\tau}}[\hat{x}]=0$. But this contradicts $F^{\mathcal{G}}_{\tilde{\tau}}[\hat{x}](d^{*},\pi^{*})<0$, in which $-d^{*}_{\tilde{\tau}}[\hat{x}]$ is the only negative term. We thus showed that if $d^{*}_{\tau}\in B_{\tau}(d^{*},\pi^{*})$ for all $\tau\in{\mathcal{T}}$, it must be that $F^{\mathcal{G}}_{\tau}[x](d^{*},\pi^{*})=0$ for all $[\tau,x]\in{\mathcal{T}}\times{\mathcal{X}}$, i.e., $d^{*}_{\tau}$ satisfies condition (SNE.1) for all $[\tau,x]\in{\mathcal{T}}\times{\mathcal{X}}$.
4 Consequences of the Reduction